OpenStudy (anonymous):

Help with matching some algebraic terms to their definitions? Mostly done, just need some intelligent guidance.

hero (hero):

Exponential Function is a function that can be expressed in the form \(f(x) = ab^x\).

hero (hero):

Base is a number raised to an exponent or logarithm

hero (hero):

The domain for exponential functions is all real numbers.

hero (hero):

Product Property of Logarithms is \(\log_{b}(xy) = \log_b(x) + \log_b(y)\).

hero (hero):

Logarithmic Function is a function that can be expressed in the form \(y = \log_{b}(x)\).

hero (hero):

\(e\) is a number derived from compound interest. Its value is 2.71828 and it is the inverse of ln.

hero (hero):

Exponential Decay is an exponential function of the form \(y = ab^{-x}\) where \(a > 0\) and \(0<b<1\)

hero (hero):

Range, for exponential growth, is \(y > 0\).

hero (hero):

Exponential Growth is an exponential function of the form \(y = ab^x\) where \(a > 0\) and \(b > 1\).

hero (hero):

Yes, for example: \(3^4\) is 3 (a number) raised to exponent 4

hero (hero):

3 is the base and 4 is the exponent.

hero (hero):

Well, this site encourages guiding the user, but sometimes, seeing things more directly helps as well. I think it helps in your case.

hero (hero):

Quotient Property = \(\log\left(\frac{x}{y}\right) = \log_b{x} - \log_b{y}\)

hero (hero):

Go ahead and post the question